OpenStudy (anonymous):

What polynomial has roots of -5, 2, and 4 ?

OpenStudy (anonymous):

the one that looks like \[(x+5)(x-2)(x-4)\] after you expand

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

Im guessing im wrong

OpenStudy (anonymous):

i would do this http://www.wolframalpha.com/input/?i=(x%2B5)(x-2)(x-4 and look at "expanded form"

OpenStudy (anonymous):

im still confused on this

OpenStudy (anonymous):

if \(-5\) is a zero, then one factor has to be \(x+5\)

OpenStudy (anonymous):

if \(2\) is a zero then one factor must be \(x-2\) and if \(4\) is a zero then one factor is \(x-4\)

OpenStudy (anonymous):

that means in "factored form" the polynomial is \[(x+5)(x-2)(x-4)\] if you want it in standard form multiply all that mess out

OpenStudy (anonymous):

well im still confused but thank you anyways

OpenStudy (anonymous):

i cannot think of another way to say it if you want to solve \((x+5)(x-2)(x-4)=0\) you would have \[x+5=0\iff x=-5\\ x-2=0\iff x=2\\ x-4=0\iff x=4\] so the zeros would be \(-5,2,4\)

OpenStudy (anonymous):

now you are told the zeros are \(-5,2,4\) so the polynomial you started with must be \[(x+5)(x-2)(x-4)\]